# Efficient Reinforcement Learning for Optimal Control with Natural Images

### v1

@@ 2.1 A Scalable Benchmark for Efficient Image Representations
We would now like to design a simple and robust benchmark for the purpose of assessing alternative image representations for reinforcement learning with natural image sequences. The number of states of the benchmark should scale to arbitrarily large numbers, its cost structure should be simple and easy to implement over long horizons, and its transition probabilities should correspond to a natural image generator that is easy to implement.
@@
The number of states of the benchmark can scale to arbitrarily large numbers by working with image patches. Image patches are arbitrarily-sized image regions extracted from a larger parent image. Discretizing the $k$th parent image into a regular grid pattern yields a set of image patches $S_{k}^{\prime}$, with each image patch denoted by $\phi_{k}(i)$; where $i\in\mathbb{Z}^{2}$ is a unique pair of coordinates giving the location of the image patch within its parent image (for example, each coordinate pair could give the coordinates of the top-left corner of an image patch). The number of image patches can be increased by decreasing the size of each image patch, similar to the case of discretizing a continuous variable where decreasing the step size increases the number of steps. The following proposition then shows the number of states of the benchmark is scalable.
@@
A state of the benchmark can be described by $i_{k}$ or $\phi_{k}(i)$, and either state is a sufficient statistic.
@@
Given any image patch $\phi_{k}(i)\in S_{k}^{\prime}$, it is possible to determine state $i\in S_{k}$ as
[…]
provided each image patch is unique. Alternatively, given state $i$, it is possible to determine state $\phi_{k}(i)$ by looking up the $i$th element of $S_{k}^{\prime}=\{\phi_{k}(1),...,\phi_{k}(n)\}$. The same is true for both $\phi_{k+1}(j)\in S_{k+1}^{\prime}$ and $j\in S_{k+1}$. Therefore, given $p_{ij}(u)$, $g_{k}(i,u,j)$, and $U_{k}(i)$; it is possible to determine $\bar{p}(\phi_{k+1}(j)|\phi_{k}(i),u)=p_{ij}(u)$, $\bar{g}_{k}(\phi_{k}(i),u,\phi_{k+1}(j))=g_{k}(i,u,j)$, and $\bar{U}_{k}(\phi_{k}(i))=U_{k}(i)$. Alternatively, given $\bar{p}(\phi_{k+1}(j)|\phi_{k}(i),u)$, $\bar{g}_{k}(\phi_{k}(i),u,\phi_{k+1}(j))$, and $\bar{U}_{k}(\phi_{k}(i))$; it is possible to determine $p_{ij}(u)=\bar{p}(\phi_{k+1}(j)|\phi_{k}(i),u)$, $g_{k}(i,u,j)=\bar{g}_{k}(\phi_{k}(i),u,\phi_{k+1}(j))$, and $U_{k}(i)=\bar{U}_{k}(\phi_{k}(i))$. But this means $i$ or $\phi_{k}(i)$ is sufficient for determining $u$ from the minimization in Eq. (4). ∎
The proof of Proposition 1 describes a very simple natural image generator $\bar{p}(\phi_{k+1}(j)|\phi_{k}(i),u)$ constructed from the transition probabilities $p_{ij}(u)$ and the image patch set $S_{k}^{\prime}$. A non-trivial model for $p_{ij}(u)$ is presented in the next sections.
The cost structure and transition probabilities of the benchmark are motivated by dragonfly tracking in videos (Loxley, 2021). The tracking task requires a tracker to follow a target as closely as possible over many time periods using a set of discrete controls that determines the dynamics of the tracker (see Fig. 1). Letting the image patch coordinates be related to the distance between the tracker and target allows this distance to be used to assign a cost to each image patch at each time period. Moving from one image patch to another changes this distance, and therefore the cost. An optimal policy corresponds to the optimal choice of controls for tracking the target starting from any image patch during any time period. This task is easily scalable to long horizons and many states provided the target dynamics scales well. Next, a scalable model for the target dynamics is presented.
2.1.1 Target Dynamics
@@
Motivated by the dragonfly video dataset used for optimal control in Loxley (2021), the following type of scenario is proposed. The target (T) starts on the bottom row of a $3\times 3$ grid (shown in Fig 1, first grid). The target and the tracker can both move up, left, or right at each time period. However, the target (but not the tracker) can also make an “evade” move by moving diagonally to evade the tracker. The tracker must therefore carefully plan its sequence of moves to follow the target closely.
During the first time period (second grid in Fig 1) the target remains where it is, while at the beginning of the second time period (third grid in Fig 1) it makes a diagonal move. An optimal tracker (blue square) anticipates a diagonal move immediately following a time period where the target is stationary, and pays a small cost to move up one square during the first time period (second grid in Fig 1). Whether the target next moves diagonally-left or diagonally-right, the optimal tracker is now able to reach the target during the next time period (third grid in Fig 1). On the other hand, a tracker applying a greedy algorithm (red squares) tries to stay as close to the target as possible during each time period, allowing the target to “evade” the greedy tracker when it makes its diagonal move, and the greedy tracker falls one move behind the target (third and fourth grids in Fig 1). A greedy tracker will always be suboptimal for this sequence.
@@
The sequence of target moves in Fig 1 can be generalized and modelled as a regular language (Sipser, 2012). The alphabet of target-moves is $\{\mathrm{s,d,r}\}$; where $\mathrm{s}\rightarrow(0,0)$, $\mathrm{d}\rightarrow(1,1)\ \text{or}\ (-1,1)$, and $\mathrm{r}\rightarrow(0,1)\ \text{or}\ (-1,0)\ \text{or}\ (1,0)$. Each symbol in the alphabet represents a change in the target coordinates, $\Delta t$, as the target moves from its current position to its next position. The regular expression given by $\mathrm{r}^{*}\cup(\mathrm{sdr})^{*}$ generates all strings in this language, representing all possible target dynamics. For example, the string given by “$\mathrm{sdr}$” represents the sequence in Fig 1: $\Delta t_{1}=(0,0)$; $\Delta t_{2}=(1,1)$; and $\Delta t_{3}=(0,1)$. According to this sequence, a target starting at $(2,1)$ stays at $(2,1)$ during the first time period; then moves to $(3,2),$ and $(3,3)$ in the second, and third time periods, respectively (see Fig 1).
A Markov chain can be constructed to generate stochastic dynamics obeying this language. To parameterize the Markov chain, we see the only degree of freedom available in the regular expression is whether a symbol of a substring begins with an “$\mathrm{r}$” or an “$\mathrm{s}$”. This can be parameterized using the probability $p$, leading to the Makov chain shown in Fig 2.
[…]
@@
Different values of $p$ give different steady-state distributions for this Markov chain. When $p=0$ or 1, the target dynamics is a deterministic repeating sequence: $(\mathrm{sdr})^{*}$ for $p=0$, and $\mathrm{r}^{*}$ for $p=1$. When $10<p<$ 1, the Markov chain is recurrent and non-periodic: i.e., when $p=0.5$ the steady-state probabilities of $\mathrm{s,d,}$ and $\mathrm{r}$ are $1/4$, $1/4$, and $1/2$, respectively; and a typical sequence looks like $\mathrm{sdrrsdrrrrsdr}$.
2.1.2 Target Tracking
It is now possible to specify the reinforcement learning benchmark in terms of a target tracking task. This task requires two dynamical variables $t_{k},c_{k}\in\mathbb{Z}^{2}$ representing the discrete two-dimensional coordinates of a target, and a controller (tracker), respectively, at time period $k$. Changes to the environment are described by the discrete dynamical system:
@@ 2.1.2 Target Tracking
A target updates its position from $t_{k}$ to $t_{k+1}$ according to the transition probabilities of the Markov chain presented in Sec. 2.1.1. In order to follow the target as closely as possible a controller updates its position from $c_{k}$ to $c_{k+1}$ by choosing a control $u_{k}$ from the set $U$ of available controls. Complete state information required for choosing optimal controls is given when the state is defined to be:
@@
where $c_{k}-t_{k}\in D$, and $\Delta t_{k-1}\in T$; so that $S_{k}=D\times T$. Here, $D,T,U\subset\mathbb{Z}^{2}$ are finite sets of integer pairs. In order to derive the transition probabilities of the corresponding controllable Markov chain, $p_{ij}(u)$, we proceed as follows. According to Eq. (11), and making use of Eqs. (9) and (10), the state update is given by
@@
and equating with Eqs. (11) and (12), then leads to $a_{1}=c_{k}-t_{k},b_{1}=\Delta t_{k-1},a_{2}=a_{1}+u_{k}-\Delta t_{k},$ and $b_{2}=\Delta t_{k}$. These equations can be viewed as constraints on the allowed values of $j$ via $a_{2}$ and $b_{2}$, and represented as $\delta(a_{2}-(a_{1}+u_{k}-\Delta t_{k}))$ and $\delta(b_{2}-\Delta t_{k})$; where $\delta(k)=1$ if $k=0$, and $\delta(k)=0$ otherwise. The controllable Markov chain is governed by the the Markov chain from Fig. 2 describing the target dynamics: $p(\Delta t_{k}|\Delta t_{k-1})$, as well as the constraints on $j$, giving:
@@
where .$\lVert.\rVert_{2}$ is the Euclidean distance. An optimal policy for the benchmark can be found after substituting these choices into Expression (4), leading to:
$\underset{u\in U}{\operatorname{min}}\ \sum_{b_{2}}p(b_{2}|b_{1})\bar{J}_{k+1}(\phi_{k+1}(a_{1}+u-b_{2},b_{2}),r_{k+1}^{*}),$ (17)
which is evaluated at each state $(a_{1},b_{1})\in D\times T$ and each time period $k$. The final choice for the benchmark is the form of the parametric function $\bar{J}_{k}$. A simple choice is a network architecture that is linear in the weights $r_{k}$ and given by the following scalar product:
$\bar{J}_{k}(\phi_{k}(i),r_{k})=r^{\top}_{k}\phi_{k}(i),$ (18)
where $\top$ is the vector transpose. The Eq. (18) makes it clear that once the weights $r_{k}$ are determined through training, $\bar{J}_{k}$ depends only on the input image $\phi_{k}$.
The benchmark is now completely specified by Eqs. (17) and (18), with Eqs. (7) and (8) used for determining $r_{k}^{*}$. The advantage of set notation is that once the sets $D,T$, and $U$ are specified, all that remains is to choose an initial state $(a_{1},b_{1})\in D\times T$ and evaluate the benchmark.
Applying the target dynamics from Sec. 2.1.1 with uniform probabilities over the “or” choices results in tracking tasks that become too uncertain over time, leading to degenerate policies that preform poorly. It is therefore necessary to work with a simpler dynamics given by $\mathrm{s}=(0,0)$, $\mathrm{d}=(1,1)$, and $\mathrm{r}=(0,1)$. The sets $T$ and $U$ then become $T=\{(0,0),(1,1),(0,1)\}$ and $U=\{(0,0),(1,0),(0,1)\}$; while the set $D$ depends on the number of coordinate differences (i.e., distances) chosen between the target and the controller. Therefore, the number of states of the benchmark is determined by choice of the size of $D$.
@@ 3 Results
Results for the reinforcement learning benchmark developed in Section 2 are now presented. We begin by looking at the optimal and greedy policies for target tracking.
@@ 3.1 Optimal and Greedy Policies in the Reinforcement-Learning Benchmark
The evasive target dynamics in the benchmark leads to different policies for optimal and greedy tracking. When $p=0$ in the Markov chain, the target dynamics is deterministic and periodic and given by the repeating sequence $(\mathrm{sdr})^{*}$. The optimal policy from Eq. (17) is given in Table 1 when the chain is initially in state $\mathrm{s}$, and cycles between three states: $i_{k}=((0,1),(0,0))$, $i_{k+1}=((0,0),(1,1))$, and $i_{k+2}=((0,0),(0,1))$.
@@
When $p=1$, the target dynamics is given by the repeating sequence $\mathrm{r}^{*}$. The optimal policy in Table 1 is now $\mu^{*}((0,0),(0,1))=(0,1)$, with cost per time period $g((0,0),(0,1))=0$. Similarly, the greedy policy in Table 2 is now $\mu^{G}((0,-1),(0,1))=(0,1)$, with cost per time period $g((0,-1),(0,1))=1$. The ratio of total costs of the greedy and optimal policies is therefore $N-1$ to 1, so the greedy policy always has a cost $N-1$ times larger than that of the optimal policy, as seen in Fig. 3 for $p=1$.
When $10<p<$ 1, the target dynamics is stochastic, recurrent, and non-periodic: at the next time period, state $\mathrm{r}$ has probability $p$ of staying in the same state, and probability $1-p$ of going to state $\mathrm{s}$. The expected total cost of the optimal policy in Table 1 now lies somewhere between the curves for the optimal policies with $p=1$ and $p=0$ in Fig. 3. Similarly, the expected total cost of the greedy policy in Table 2 now lies between the curves for the greedy policies with $p=0$ and $p=1$ in Fig. 3. The optimal policy and its expected total cost were found using both fitted value iteration and the dynamic programming algorithm, with both approaches yielding the same results. Results for the greedy policy were found directly from Eq. (1) with $\mu_{k}(i)=\mu_{k}^{G}(i)$ from Eq. (19) using an efficient algorithm described in the Appendix. These results were also confirmed using policy evaluation; that is, solving $J_{k}(i)=\sum_{j}p_{ij}(\mu_{k}^{G}(i))(g_{k}(i,\mu_{k}^{G}(i),j)+J_{k+1}(j))$ using dynamic programming iterations. In this case, $J_{N}(i)$ was set to the terminal cost as for fitted value iteration and dynamic programming.

### v2

@@ 3 A Scalable Benchmark for Efficient Image Representations
We would now like to design a simple and robust benchmark for the purpose of assessing alternative image representations for optimal control with natural image sequences. The number of states of the benchmark should scale to arbitrarily large numbers, and its cost structure and dynamics should be easy to implement over long horizons.
@@
The number of states of the benchmark can scale to arbitrarily large numbers by working with image patches. An image patch is an arbitrarily sized image region extracted from a larger parent image. Discretizing a parent image into a regular grid pattern yields a set of image patches $\bar{S}$, with each image patch $\phi(i)\in\bar{S}$ indexed by $i\in\mathbb{Z}^{2}$: a unique pair of coordinates giving the location of the image patch within its parent image (for example, each coordinate pair could give the coordinates of the top-left corner of an image patch). Given a parent image of fixed size the number of image patches can be increased simply by decreasing the size of each image patch. The following proposition will be used to help construct the benchmark.
@@
A state of the benchmark can be described by $i$ or $\phi(i)$, and either state is a sufficient statistic.
@@
Given any image patch $\phi(i)\in\bar{S}$, it is possible to determine its index $i$ by running through all image patches to find the closest match:
$\hat{i}=\underset{m}{\operatorname{argmin}}\lVert\phi(m)-\phi(i)\rVert_{2}^{2},$
provided each image patch is unique. This index also corresponds to the state $i\in S$. Alternatively, given state $i$, it is possible to determine the image patch $\phi(i)$ by looking up the $i$th element of $\bar{S}=\{\phi(1),...,\phi(n)\}$. The same is true for both $j$ and $\phi(j)$. Therefore, given $p_{ij}(u)$, $g_{k}(i,u,j)$, and $U_{k}(i)$; it is possible to determine $\bar{p}(\phi(j)|\phi(i),u)=p_{ij}(u)$, $\bar{g}_{k}(\phi(i),u,\phi(j))=g_{k}(i,u,j)$, and $\bar{U}_{k}(\phi(i))=U_{k}(i)$. Alternatively, given $\bar{p}(\phi(j)|\phi(i),u)$, $\bar{g}_{k}(\phi(i),u,\phi(j))$, and $\bar{U}_{k}(\phi(i))$; it is possible to determine $p_{ij}(u)=\bar{p}(\phi(j)|\phi(i),u)$, $g_{k}(i,u,j)=\bar{g}_{k}(\phi(i),u,\phi(j))$, and $U_{k}(i)=\bar{U}_{k}(\phi(i))$. ∎
The proof of Proposition 1 describes a very simple image generator $\phi(j)\sim\bar{p}(\boldsymbol{\cdot}|\phi(i),u)$ constructed from an image patch set $\bar{S}$, and any set of transition probabilities $p_{ij}(u)$. A non-trivial model for $p_{ij}(u)$ is presented in the next sections.
The cost structure and transition probabilities of the benchmark are motivated by dragonfly tracking in videos (loxley2). The tracking task requires a tracker to follow a target as closely as possible over many time periods using a set of discrete controls that determines the dynamics of the tracker (see Figure 2). Letting the image patch coordinates be related to the distance between the tracker and target allows this distance to be used to assign a cost to each image patch. Moving from one image patch to another changes this distance, and therefore the cost. An optimal policy corresponds to the optimal choice of controls for tracking the target starting from any image patch during any time period. This task is easily scalable to long horizons and many states provided the target dynamics scales well. Next, a scalable model for the target dynamics is presented.
3.1 Target Dynamics
@@
Motivated by the dragonfly video dataset used for optimal control in loxley2, the following type of scenario is proposed. The target (T) starts on the bottom row of a $3\times 3$ grid (shown in Fig 1, first grid). The target and the tracker can both move up, left, or right at each time period. However, the target (but not the tracker) can also make an “evade” move by moving diagonally to evade the tracker. The tracker must therefore carefully plan its sequence of moves in order to follow the target closely.
During the first time period (second grid in Fig 1) the target remains where it is, while at the beginning of the second time period (third grid in Fig 1) it makes a diagonal move. An optimal tracker (blue square) anticipates a diagonal move immediately following a time period where the target is stationary, and pays a small cost to move up one square away from the target during the first time period (second grid in Fig 1). Whether the target next moves diagonally left or diagonally right, the optimal tracker is now able to reach the target during the next time period (third grid in Fig 1). On the other hand, a tracker applying a greedy algorithm (red squares) tries to stay as close to the target as possible during each time period, allowing the target to “evade” the greedy tracker when it makes its diagonal move, causing the greedy tracker to fall one move behind the target (third and fourth grids in Fig 1). A greedy tracker will always be suboptimal for this sequence.
@@
The sequence of target moves in Fig 1 can be generalized and modelled as a regular language (sipser). The alphabet of target-moves is $\{\mathrm{s,d,r}\}$; where $\mathrm{s}\rightarrow(0,0)$, $\mathrm{d}\rightarrow(1,1)\ \text{or}\ (-1,1)$, and $\mathrm{r}\rightarrow(0,1)\ \text{or}\ (-1,0)\ \text{or}\ (1,0)$. Each symbol in the alphabet represents a change in the target coordinates, $\Delta t$, as the target moves from its current position to its next position. The regular expression given by $\mathrm{r}^{*}\cup(\mathrm{sdr})^{*}$ generates all strings in this language, representing all possible target dynamics. For example, the string given by $\mathrm{sdr}$” represents the sequence in Fig 1: $\Delta t_{1}=(0,0)$; $\Delta t_{2}=(1,1)$; and $\Delta t_{3}=(0,1)$. According to this sequence, a target starting at $(2,1)$ stays at $(2,1)$ during the first time period; then moves to $(3,2),$ and $(3,3)$ in the second, and third time periods, respectively (see Fig 1).
A Markov chain can be constructed to generate stochastic dynamics obeying this language. The Markov chain is parameterized by the single degree of freedom available in the regular expression: whether a symbol of a substring begins with an $\mathrm{r}$” or an $\mathrm{s}$”. This is parameterized using the probability $p$, leading to the Makov chain shown in Fig 2.
rppd11s1p1-p11
@@
Different values of $p$ give different steady state distributions of the Markov chain. When $p=0$ or 1, the target dynamics is a deterministic periodic sequence: $(\mathrm{sdr})^{*}$ for $p=0$, and $\mathrm{r}^{*}$ for $p=1$. When $0<p<1$, the Markov chain is recurrent and non-periodic: i.e., when $p=0.5$ the steady state probabilities of $\mathrm{s,d,}$ and $\mathrm{r}$ are $1/4$, $1/4$, and $1/2$, respectively; and a typical sequence of twelve target moves looks like $\mathrm{sdrrsdrrrrsdr}$.
3.2 Target Tracking
It is now possible to specify the reinforcement learning benchmark as a target tracking task. This task requires two dynamical variables $t_{k},c_{k}\in\mathbb{Z}^{2}$ representing the discrete two-dimensional coordinates of a target, and a controller (tracker), respectively, at time period $k$. Changes to the environment are described by the discrete dynamical system:
@@ 3.2 Target Tracking
A target updates its position from $t_{k}$ to $t_{k+1}$ according to the transition probabilities of the Markov chain presented in Sec. 3.1. In order to follow the target as closely as possible a controller updates its position from $c_{k}$ to $c_{k+1}$ by choosing a control $u_{k}$ from the set $U$ of available controls. The state information required for choosing an optimal control is given by defining the state to be:
@@
where $c_{k}-t_{k}\in D$, and $\Delta t_{k-1}\in T$; so that $S=D\times T$. Here, $D,T,U\subset\mathbb{Z}^{2}$ are sets of integer pairs. In order to derive the transition probabilities $p_{ij}(u)$, we proceed as follows. According to Eq. (11), and making use of Eqs. (9) and (10), the state update is given by
@@
and upon equating with Eqs. (11) and (12), leads to $a_{1}=c_{k}-t_{k},b_{1}=\Delta t_{k-1},a_{2}=a_{1}+u_{k}-\Delta t_{k},$ and $b_{2}=\Delta t_{k}$. These equations can be viewed as constraints on the allowed values of $j$ via $a_{2}$ and $b_{2}$, and represented as $\delta(a_{2}-(a_{1}+u_{k}-\Delta t_{k}))$ and $\delta(b_{2}-\Delta t_{k})$; where $\delta(k)=1$ if $k=0$, and $\delta(k)=0$ otherwise. The controllable Markov chain is governed by the the Markov chain from Fig. 3 describing the target dynamics: $p(\Delta t_{k}|\Delta t_{k-1})$, as well as the constraints on $j$, giving:
@@
where $\lVert.\rVert_{2}$ is the Euclidean distance. Substituting these choices into Expression (3), then making use of Proposition 1 to replace $J_{k+1}^{*}(j)$ with $\bar{J}_{k+1}(\phi(j),r_{k+1}^{*})$, leads to:
$\underset{u\in U}{\operatorname{min}}\ \sum_{b_{2}}p(b_{2}|b_{1})\bar{J}_{k+1}(\phi(a_{1}+u-b_{2},b_{2}),r_{k+1}^{*}).$ (17)
The expressions (7) and (8) are used for determining $r_{k}^{*}$ with (6) used for $\bar{J}_{k}$. The sets $T$ and $U$ are specified by $T=\{(0,0),(1,1),(0,1)\}$ and $U=\{(0,0),(1,0),(0,1)\}$; restricting the target moves as in $T$ leads to a tracking task with less uncertainty and better optimal policies. The set $D$ determines the number of coordinate differences (i.e., distances) between the target and the controller, and depends on the number of states (image patches) chosen for the benchmark. Once $D$ has been specified, an optimal policy for the benchmark is found by carrying out the minimization in (17) at each state $(a_{1},b_{1})\in D\times T$ and at each time period $k$.
@@ 4 Results
Results for the reinforcement learning benchmark developed in Section 3 are now presented. All code for the following figures is supplied at Loxley4. We begin by looking at the optimal and greedy policies for target tracking.
@@ 4.1 Optimal and Greedy Policies in the Reinforcement-Learning Benchmark
Evasive target dynamics in the benchmark leads to different policies for optimal and greedy tracking. When $p=0$ in the Markov chain, the target dynamics is deterministic and periodic and given by the repeating sequence $(\mathrm{sdr})^{*}$. The optimal policy from Eq. (17) is given in Table 1 when the chain is initially in state $\mathrm{s}$, and cycles between three states: $i_{k}=((0,1),(0,0))$, $i_{k+1}=((0,0),(1,1))$, and $i_{k+2}=((0,0),(0,1))$. Remember, the first pair of each state is the coordinate difference between the target and controller (tracker), while the second pair is the target’s change in coordinates during the previous time period.
@@
When $p=1$, the target dynamics is given by the repeating sequence $\mathrm{r}^{*}$. The optimal policy in Table 1 is now $\mu^{*}((0,0),(0,1))=(0,1)$, with cost per time period $g((0,0),(0,1))=0$. Similarly, the greedy policy in Table 2 is now $\mu^{G}((0,-1),(0,1))=(0,1)$, with cost per time period g((0,1),(0,1))=1g((0,-1),(0,1))=1. The ratio of total costs of the greedy to optimal policies is therefore N1N-1 to 1, so the greedy policy always has a cost N1N-1 times larger than that of the optimal policy, as seen in Fig. 4 for $p=1$.
When $0<p<1$, the target dynamics is stochastic, recurrent, and non-periodic: at the next time period, state $\mathrm{r}$ has probability $p$ of staying in the same state, and probability 1p1-p of going to state $\mathrm{s}$. The expected total cost of the optimal policy in Table 1 now lies somewhere (not shown) between the curves for the optimal policies with $p=1$ and $p=0$ in Fig. 4. Similarly, the expected total cost of the greedy policy in Table 2 now lies (not shown) between the curves for the greedy policies with $p=0$ and $p=1$ in Fig. 4.
The optimal policy and its expected total cost were found using both fitted value iteration and the dynamic programming algorithm, with both approaches yielding the same result. Results for the greedy policy were found directly from Eq. (1) with $\mu_{k}(i)=\mu_{k}^{G}(i)$ from Eq. (18) using an efficient algorithm described in the Appendix. These results were also confirmed using policy evaluation; that is, by solving $J_{k}(i)=\sum_{j}p_{ij}(\mu_{k}^{G}(i))(g_{k}(i,\mu_{k}^{G}(i),j)+J_{k+1}(j))$ using dynamic programming iterations. In this case, $J_{N}(i)$ was set to the terminal cost as for fitted value iteration and dynamic programming.